# FOSS: A Self-Learned Doctor for Query Optimizer

### v1

@@ III Planner
Algorithm 1 shows the overall training process of planner. When a query $Q$ of a workload $W$ is inputted, the DBMS optimizer generates the original plan $CP_{0}$ (Line 2). From $CP_{0}$, the planner extracts an incomplete plan $ICP_{0}$ composed of the join order and join methods (Line 3). The planner initializes the history plan set $T$ and the estimated best plan $CP_{best}$ (Line 4-5). For every step, the planner calculates the forbidden actions based on the $ICP_{t-1}$ (Line 7). If the previous action is $Swap$ operator, the current action will be restricted to only changing the adjacent join method (Line 8-10). Next, the planner represents the previous state based on the $CP_{t-1}$ and step status (Line 11-12) and evaluates the action based on the state representation $statevec_{t-1}$ and $actionmask$ (Line 13). Then it applies this action to $ICP_{t-1}$ and produces a new incomplete plan $ICP_{t}$ (Line 14). The $ICP_{t}$ will be completed by the DBMS, resulting in a complete plan $CP_{t}$ (Line 15). The planner computes the penalty value and uses it to initialize the reward value (Line 16). If the current incomplete plan $ICP_{t}$ has not appeared in this episode, we add the bounty to reward and add the $ICP_{t}$ to $T$ (Line 17-20). If the current plan is better than the optimal plan, the planner updates the estimated optimal plan $CP_{best}$ (Line 21). Experiences are collected in order to update the agent (Line 22). Finally, upon concluding the exploration of the current workload, the agent is updated using $experiences$ (Line 25).
@@ IV Asymmetric Advantage Model
In this section, we will provide a detailed description of the asymmetric advantage model. We begin by introducing the state representation, followed by the advantage model. Finally, we will discuss the design of the loss function.
@@ IV-B Advantage Model
As described in (3), we need to calculate the advantage between two plans. In the real environment, precise advantage values can be obtained by executing the plans using DBMS executor. In simulated environment, we estimate the advantage values using an advantage model which takes plan pairs $(CP_{l},CP_{r})$ as input. It initially encodes the two input plans separately, transforming them into $(statevec_{l},statevec_{r})$ using the state network. Next, it incorporates position information for each $statevec$ to obtain $(statevec_{l}\oplus pos_{left},statevec_{r}\oplus pos_{right})$. Finally, the pair is passed to the final output layer, where it’s mapped to the scores (i.e., the output of $Adv$). The asymmetric advantage model $\theta_{adv}$ can be described as follows:
@@
Considering that the dynamic changes in system workload may cause the labels of plan pairs with similar execution latency to change continuously and to select significantly superior plans, we take the point set $\{0.05,0.75\}$ to divide the interval $(-\infty,1]$, resulting in three scores (i.e., $\{0,1,2\}$) for the output of the $Adv$.
@@ IV-C Asymmetric Loss
Since plans obtained from traditional optimizer are often of reasonable performance, starting with such plans as baselines and making modifications often yields more inferior plans. As a result, training samples for $\theta_{adv}$ are skewed towards a larger number of samples with a score of 0 and fewer samples with a score of 2. This uneven distribution of samples across different labels will adversely affect the training of the $\theta_{adv}$. Inspired by Asymmetric Loss [17], FOSS addresses the issue by increasing the weights of difficult-to-classify samples. Additionally, to prevent overfitting during training, FOSS also employs label smoothing.
@@ V Simulated Learner
In this section, we will begin by introducing the simulated environment. Next, we will provide an explanation of how FOSS explores the real environment and trains the agent using the simulated environment.
@@ V-A Simulated Environment
As described in Algorithm 1, data in the form of $\{s,a,r,s^{{}^{\prime}}\}$ are collected to update the agent. However, in many environments, obtaining the real state and reward is challenging due to high costs or security concerns. To overcome this challenge, researchers propose simulating environment using neural networks to accelerate agent’s learning. An simulated environment always includes State Transition Dynamics Model $\theta_{p}(s,a)\to s^{{}^{\prime}}$ and Reward Function Model $\theta_{r}(s,a)\to r$ [18]. They will have the agent interact with the real environment and then use the real experiences to perform supervised training of the simulated environment [19]. Then the agent efficiently interacts with the simulated environment to generate simulated experiences, which are then used to update the agent accordingly.
In the context of planner, state is consist of complete plan $CP$ and step status, while step status is determined and $CP$ can be obtained from DBMS optimizer with low cost, especially within the guidance of $ICP$. Therefore, we can assign $\theta_{p}(s,a)$ with DBMS optimizer $\Gamma_{p}$. However, obtaining the true reward requires executing the complete plan, which can be time-consuming. We only need to be aware of how we can simulate $\theta_{r}(s,a)$ to provide rewards feedback timely. Using the asymmetric advantage model $\theta_{adv}$ that we have established to assign rewards is a natural idea. Therefore, we employ the DBMS optimizer $\Gamma_{p}$ and $\theta_{adv}$ to construct a simulated environment $\hat{E}(\Gamma_{p},\theta_{adv})$.
@@ V-B Training Loop
In this subsection, we will discuss how to train the agent of the planner effectively using the simulated environment and the real environment, and how to collect training data for updating $\theta_{adv}$. It can be divided into three stages: the exploration phase, the training phase and the validation phase.
Exploration Phase. For a batch of training workload queries, the agent interacts with real environment to perform plan exploration and collect the execution latency into execution buffer. However, during the initial stage of exploration, agent is possible to generate catastrophic plans that are significantly worse than the original plan, and it may even lead to timeouts, significantly slowing down the exploration phase. To address this issue, we introduce dynamic timeouts to terminate the execution of poorly performing plans. The timeout is set to $1.5$ times the original plan’s execution latency. If the execution latency of a new generated plan exceeds this threshold, it is terminated and labeled as timeout. Then, as described in IV-B, we utilize the data in execution buffer to form the training data for $\theta_{adv}$ and update it. We will filter out plan pairs that involve timeouts for both. Although the timeout strategy results in a loss of training samples, it significantly increases the data collection rate.
Training Phase. FOSS randomly samples from the history queries and acquires simulated experiences by interacting with the simulated environment $\hat{E}(\Gamma_{p},\theta_{adv})$, following the process outlined in the Algorithm 1. Meanwhile, FOSS will collect excellent plans based on the outcomes of $\theta_{adv}$.
Validation Phase. FOSS executes the plans collected during the training phase, records their execution latency into the execution buffer, and updates $\theta_{adv}$ accordingly. This process enables the timely rectification of errors in $\theta_{adv}$, preventing the accumulation of errors in DRL training due to inaccuracies in reward estimation.
Global Process. FOSS initiates an exploration phase, succeeded by multiple rounds of training phases and a validation phase after each training iteration. This cyclic process is repeated several times until the FOSS converges. The three stages can be conducted interactively or simultaneously.
@@ VI-A Experiment Setup
Experiment Support. We deploy FOSS alongside the SOTA methods on an Ubuntu 20.04 system, featuring a hardware configuration that includes a 2.40GHz Intel Xeon E5-2640 CPU, 256GB of memory, and an NVIDIA Tesla P40 GPU. We implement FOSS using PyTorch and utilize Ray [20] for the RL component. We utilize PPO [21] as the base RL algorithm due to its effectiveness in mitigating significant differences in the action distribution before and after agent updates through KL divergence. It further ensures the accuracy of the estimated reward provided by $\theta_{adv}$ in the simulated environment.
@@
Comparision. We employ PostgreSQL as baseline and compare FOSS with the SOTA methods, including Bao [12], Balsa [8], and HybridQO [14].
@@
Bao follows plan-steerer framework. It steers traditional optimizer to select better query plans by employing a value model to choose different arms (i.e., hint sets). By default configuration, it has 5 arms available. We configure it according to the instructions provided in their documents [23].
Balsa is an end-to-end query optimizer that operates independently from traditional optimizer, following plan-constructor framework. We follow the instructions outlined in its source code [24] to deploy it .
HybridQO combines learned-based optimizer with traditional optimizer. It also follows plan-steerer framework. It employs Monte Carlo Tree Search to discover promising leading join order and uses these as hints to guide traditional optimizer in generating better query plans. We replicate their experimental setup as described in [25].
@@
Metrics. We deploy all learned query optimizers on the training workload and evaluate their performance on the testing workload using the two metrics: i) Workload Relevant Latency (WRL): it primarily reflects the overall latency level of the entire workload, as previously employed in the approach [13, 8, 9]. Its computation involves the ratio of the total latency of the entire workload executed by the learned-based optimizer to the total latency observed in expert optimizer. ii) Geometric Mean Relevant Latency (GMRL): it is introduced in [6] and demonstrates the optimization effectiveness at the query level. It can be expressed using the formula $GMRL=\sqrt[|W|]{\prod_{q\in W}\frac{ET_{learned}(q)}{ET_{expert}(q)}}$, where $ET_{learned}$ and $ET_{expert}$ respectively represent the execution latency of the learned-based optimizer and the expert optimizer. For the two metrics, a smaller value indicates better performance. A value greater than 1 implies inferiority compared to the expert optimizer, while a value less than 1 signifies the opposite.
@@ VI-B FOSS Performance and Comparison
Unless specified otherwise, we set the $maxsteps$ to 3, explore 10 queries in each round of the exploration phase, and use 900 episodes (i.e., 900 queries) for each update of agent during training phase. We deploy all methods on each benchmark, running them five times with different random seeds.
(a) Training workload
(b) Testing workload
@@
VI-B1 FOSS Performance Overview
(a) WRL of JOB
(b) WRL of TPC-DS
(c) WRL of STACK
(d) GMRL of JOB
(e) GMRL of TPC-DS
(f) GMRL of STACK
Figure 4: The training curves of FOSS on various workloads.
In this subsection, we analyze the training efficiency and performance of FOSS. Overall, FOSS outperforms the expert optimizer (i.e., PostgreSQL optimizer) with speedup of 5.88×, 1.28×, and 1.37× in total latency for the training workload (JOB, TPC-DS, STACK) and 2.78×, 1.30×, and 1.18× for the testing workload, respectively (Fig. 3).
Fig. 4 illustrates the training curves of FOSS on JOB, TPC-DS, and STACK. The solid red line represents the average change in metrics over five runs on the training workload, while the blue line indicates the average change in metrics on the testing workload. The shaded area indicates the range between the minimum and maximum values. The line at 1.0 indicates the performance of the expert optimizer. Benefiting from the assurance provided by the original plans, FOSS can swiftly demonstrate its optimization effects and surpass the performance of expert optimizer.
On the JOB workload, the complexity of queries and data distribution allows FOSS to have more room for optimization. After about 8 hours of training, FOSS achieves convergence, ultimately saving $64\%$ of the latency on the testing workload relative to expert optimizer. A steady decrease on the GMRL curve until convergence to $0.64$ on the testing workload demonstrates FOSS’s ability to optimize the majority of original plans and achieve high performance. On the training workload, FOSS shows a similar convergence trend, indicating its good generalization.
On the TPC-DS and STACK workloads, FOSS occasionally encounters catastrophic plans during the training process, leading to sudden increases in WRL and GMRL. However, it can avoid these plans as training progresses, eventually converging and achieving relatively good results. Due to their simpler workload, their training processes converge in 1-3 hours and have low tail variance.
VI-B2 Performance Comparison
@@
Methods WRL, train GMRL, train WRL, test GMRL, test Training time (hours)
@@
PostgreSQL 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 / / /
Bao 0.22 0.78 1.58 0.54 0.96 1.18 0.44 0.78 1.37 0.84 0.95 1.78 3.35 1.32 3.03
Balsa 0.19 1.79 NA 0.62 1.98 NA 0.45 1.86 NA 1.25 2.08 NA 12.18 10.58 TLE
HybridQO 0.61 0.95 1.43 0.71 1.07 1.40 0.51 0.92 0.86 0.75 1.04 1.01 3.51 0.52 1.58
FOSS 0.17 0.78 0.73 0.48 0.87 0.78 0.36 0.77 0.85 0.64 0.87 0.87 8.09 1.55 2.31
@@
In this subsection, we compare the performance of FOSS with the SOTA methods. We ensure that all methods are evaluated after convergence. Notably, Balsa’s training on STACK is hindered by catastrophic plans generated during the initial phase, preventing it from completing within a reasonable timeframe.
As described in Fig. 3, despite the left-deep restriction, FOSS consistently outperforms or matches the other three methods for both training and testing workloads. FOSS outperforms Bao, HybridQO, and Balsa with average speedups of 1.48×, 2.26×, and 1.71× on training workloads, and 1.28×, 1.21×, and 1.84× on testing workloads, respectively.
As shown in Table I, FOSS not only outperforms other methods at the workload total latency but also exhibits superior performance in GMRL metrics. Worth mentioning is the performance on the JOB workload, where Bao and Balsa exhibit a relatively small difference compared to FOSS in WRL, but a more significant gap emerges in GMRL. This discrepancy arises from FOSS efficiently interacting with the simulated environment, allowing for more exploration compared to Balsa. Compared to Bao, FOSS employs finer-grained optimization, resulting in a higher likelihood of generating better plans. These factors enable FOSS to surpass the expert optimizer across a broader range of queries, indicating its stronger optimization capabilities. This will be further confirmed in VI-B4. However, these advantages come with an associated cost, leading to relatively longer training time on the JOB compared to HybridQO and Bao. Nevertheless, in comparison to Balsa, another MDP-based method, FOSS exhibits a significant reduction in training time. While on STACK and TPC-DS, the training time difference among FOSS, Bao, and HybridQO is not pronounced and is much less than that of Balsa, FOSS surpasses all other methods, demonstrating the superiority of our approach.
VI-B3 Optimization Time Comparison
Figure 5: Optimization time of various optimizers.

### v2

@@ III Planner
Algorithm 1 shows the overall training process of planner. When a query $Q$ is inputted, the DBMS optimizer generates the original plan $CP^{Q}_{ORI}$ (Line 2). From $CP^{Q}_{ORI}$, the planner extracts the $ICP^{e}_{0}$ composed of the join order and join methods (Line 3). The planner initializes the episode buffer set $T$ and the estimated optimal plan $\overline{CP}^{e}_{t}$ (Line 4-5). For every step, the planner calculates the forbidden actions based on the $ICP^{e}_{t-1}$ (Line 7). If the previous action is $Swap$ operator, the current action will be restricted to only overriding the adjacent join method (Line 8-10). Next, the planner represents the previous state based on the $CP^{e}_{t-1}$ and step status (Line 11-12) and evaluates the action based on the state representation $statevec^{e}_{t-1}$ and $actionmask$ (Line 13). Then it applies this action to $ICP^{e}_{t-1}$ and produces a new incomplete plan $ICP^{e}_{t}$ (Line 14). The $ICP^{e}_{t}$ will steer the DBMS optimizer to output a complete plan $CP^{e}_{t}$ (Line 15). The planner computes the penalty value and uses it to initialize the reward (Line 16). If the current incomplete plan $ICP^{e}_{t}$ has not appeared in this episode, we add the bounty to reward and add the $ICP^{e}_{t}$ to $T$ (Line 17-20). If the current plan is better than $\overline{CP}^{e}_{t-1}$, the planner updates the $\overline{CP}^{e}_{t}$ with $CP^{e}_{t}$ (Line 21). Experiences are collected in order to update the agent (Line 22). Finally, upon concluding the exploration of the current workload, the agent is updated using $experiences$ (Line 25).
@@ IV Asymmetric Advantage Model
In this section, we will provide a detailed description of the AAM (i.e., asymmetric advantage model). We begin by introducing the state representation, followed by the advantage model. Finally, we will discuss the design of the loss function.
@@ IV-B Advantage Model
As described in (2), we need to calculate the advantage between two plans. In the real environment, precise advantage values are obtained by executing the plans offline. In simulated environment, we estimate the advantage values using an advantage model which takes plan pairs $(CP_{l},CP_{r})$ as input. It initially encodes the two input plans separately, transforming them into $(statevec_{l},statevec_{r})$ using the state network. Next, it incorporates position information for each $statevec$ to obtain $(statevec_{l}\oplus pos_{left},statevec_{r}\oplus pos_{right})$. Finally, the pair is passed to the final output layer, where it’s mapped to the scores (i.e., the output of $Adv$). The AAM $\theta_{adv}$ can be described as follows:
@@
Considering that the dynamic changes in system workload may cause the labels of plan pairs with similar execution latency to change continuously and to select significantly superior plans, we take the point set $\{0.05,0.50\}$ to divide the interval $(-\infty,1]$, resulting in three scores (i.e., $\{0,1,2\}$) for the output of the $Adv$.
@@ IV-C Asymmetric Loss
Since plans obtained from traditional optimizer are often of reasonable performance, starting with such plans as baselines and making modifications often yields more inferior plans. As a result, training samples for $\theta_{adv}$ are skewed towards a larger number of samples with a score of 0. This uneven distribution of samples across different labels will adversely affect the training of the $\theta_{adv}$. Inspired by Asymmetric Loss [20], FOSS addresses the issue by increasing the weights of difficult-to-classify samples. Additionally, to prevent overfitting during training, FOSS also employs label smoothing.
@@ V Simulated Learner
In this section, we will begin by introducing the simulated environment. Next, we will provide an explanation of how FOSS explores the real environment and trains the agent with the simulated environment.
@@ V-A Simulated Environment
As described in Algorithm 1, data in the form of $\{s,a,r,s^{{}^{\prime}}\}$ are collected to update the agent. However, in many environments, obtaining the real state and reward is challenging due to high costs or security concerns. To overcome this challenge, researchers propose simulating environment using neural networks to accelerate agent’s learning. A simulated environment always includes State Transition Dynamics Model $\theta_{p}(s,a)\to s^{{}^{\prime}}$ and Reward Function Model $\theta_{r}(s,a)\to r$ [21]. They will have the agent interact with the real environment and use the real experiences to perform supervised training of the simulated environment [22]. Then the agent efficiently interacts with the simulated environment to generate simulated experiences, which are used to update the agent accordingly.
In the context of planner, state is consist of $CP$ and step status, while step status is determined and $CP$ can be obtained from DBMS optimizer with low cost, especially within the guidance of $ICP$. Therefore, we can assign $\theta_{p}(s,a)$ with DBMS optimizer $\Gamma_{p}$. However, obtaining the true reward requires executing the complete plan, which can be time-consuming. We only need to be aware of how we can simulate $\theta_{r}(s,a)$ to provide rewards feedback timely. Using the AAM $\theta_{adv}$ that we have established to assign rewards is a natural idea. Therefore, we employ the DBMS optimizer $\Gamma_{p}$ and $\theta_{adv}$ to construct a simulated environment $\hat{E}(\Gamma_{p},\theta_{adv})$.
@@ V-B Training Loop
Figure 3: The training loop of FOSS.
In this subsection, we will discuss how to train the agent of the planner effectively using the simulated environment and the real environment, and how to collect training data for updating AAM.
As depicted in Fig. 3, FOSS concurrently conducts training for the planner, AAM, and plan execution. It begins by randomly initializing the planner, which then generates candidate plans for sampled queries from the workload. After executing these plans, FOSS collects their execution latency into the execution buffer. Then, as described in IV-B, we utilize the data in execution buffer to form the training data for AAM and update it. Once AAM training concludes, the planner interacts with the simulated environment $\hat{E}(\Gamma_{p},\theta_{adv})$, following Algorithm 1. Simultaneously, FOSS gathers promising plans to be executed based on the feedback of AAM during planner training. FOSS also periodically samples queries from the workload at random and collects candidate plans to be executed. The listener process triggers plan execution by the executor upon receiving new plans to be executed and aggregates execution latency into the execution buffer. Upon accumulating a sufficient number of executed plans, FOSS reorganizes the training data from the execution buffer and retrains the AAM. During plan execution, we implement dynamic timeouts to prevent poorly performing plans from significantly hindering the training progress. The timeout duration is set to 1.5 times the execution latency of the original plan. If the execution latency of a newly generated plan exceeds this threshold, it is terminated and labeled as a timeout. When organizing the AAM training data, we filter out plan pairs that involve timeouts for both plans. While this timeout strategy may lead to a reduction in the number of training samples for AAM and potentially lower the prediction accuracy of AAM, it significantly increases the rate of data collection and improves the training efficiency of FOSS.
@@ VI-A Experiment Setup
Experiment Support. We deploy FOSS alongside the SOTA methods on an Ubuntu 18.04 system, featuring a hardware configuration that includes Intel(R) Xeon(R) Gold 5118 CPU @2.30GHz, 256GB of memory, and a Geforce RTX 3090 GPU. We implement FOSS using PyTorch and utilize Ray [23] for the RL component. We utilize PPO [24] as the base RL algorithm due to its effectiveness in mitigating differences in the action distribution before and after agent updates through KL divergence. It further ensures the accuracy of the estimated reward provided by AAM in the simulated environment.
@@
Comparision. We employ PostgreSQL as baseline and compare FOSS with the SOTA methods, including the two plan-steerer methods (i.e., Bao [13] and HybridQO [15]), and the two plan-constructor methods (i.e., Balsa [9] and Loger [10]).
@@
Bao steers traditional optimizer to select better plan by employing a value model to choose suitable hint set. By default configuration, it has 5 hint sets available. We configure it according to the instructions [26].
HybridQO combines learned-based optimizer with traditional optimizer. It employs Monte Carlo Tree Search to discover promising leading join order and uses these as hints to steer traditional optimizer to generate better plans. We replicate their experimental setup as described in [27].
Balsa is an end-to-end query optimizer that operates independently from traditional optimizer. We follow the instructions outlined in its source code [28] to deploy it.
Loger focuses on join order and join method optimization, which aligns with FOSS. It learns to join individual tables in a bottom-up manner but restricts specific join methods instead of directly selecting one for each join. We implement it according to its source code [29].
@@
Metrics. We deploy all learned query optimizers on the training workload and evaluate their performance on the testing workload using the two metrics: i) Geometric Mean Relevant Latency (GMRL): it is introduced in [7] and demonstrates the optimization effectiveness at the query level. It can be expressed using the formula $GMRL=\sqrt[|W|]{\prod_{q\in W}\frac{ET_{l}^{q}}{ET_{e}^{q}}}$, where $ET_{l}$ and $ET_{e}$ respectively represent the execution latency of the learned-based optimizer and the expert optimizer. ii) Workload Relevant Latency (WRL): it primarily reflects the overall latency level of the entire workload, as previously employed in the approach [14, 9, 10]. It can be expressed using $WRL=\frac{\sum_{q}^{W}(ET^{q}_{l}+OT^{q}_{l})}{\sum_{q}^{W}(ET^{q}_{e}+OT^{q}_{e})}$, where $OT_{l}$ and $OT_{e}$ respectively represent the optimization time of the learned-based optimizer and the expert optimizer. For the two metrics, a smaller value indicates better performance. A value greater than 1 implies inferiority compared to the expert optimizer, while a value less than 1 signifies the opposite.
@@ VI-B FOSS Performance and Comparison
Unless specified otherwise, we set the $maxsteps$ to 3 and use 900 episodes for each update of agent during the training phase of planner. We deploy all methods on each benchmark, running them three times with different random seeds.
@@
VI-B1 Performance Overview
@@
Methods WRL / train GMRL / train WRL / test GMRL / test Workload Runtime (s)
@@
PostgreSQL 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 161.50 35.59 36.31
Bao 0.28 0.86 1.42 0.70 0.92 0.81 0.19 0.88 1.50 0.71 0.96 0.95 30.49 31.32 54.47
Balsa 0.30 1.66 TLE 1.30 2.54 TLE 0.31 1.72 TLE 1.69 2.59 TLE 49.76 61.22 TLE
Loger 0.20 1.20 0.57 0.60 1.06 0.64 0.18 1.22 0.82 0.64 1.09 0.77 29.07 42.71 29.77
HybridQO 0.36 0.98 1.56 0.65 0.95 0.92 0.23 0.98 1.57 0.74 1.04 0.97 36.92 34.88 57.01
FOSS 0.16 0.87 0.57 0.48 0.90 0.63 0.12 0.87 0.80 0.61 0.91 0.72 19.38 30.96 29.05
Figure 5: Training curves of speedup relative to expert optimizer.
@@
In this subsection, we compare the overall performance of FOSS with the other optimizers. We ensure that all SOTA methods are evaluated after convergence. Table I shows the results of each method on various workloads. Notably, Balsa’s training on Stack is hindered by catastrophic plans generated during the initial phase, preventing it from completing within a reasonable timeframe. Fig. 4 illustrates the relative speedup of FOSS compared to other optimizers across various workloads.
Overall, FOSS outperforms the expert optimizer (i.e., PostgreSQL optimizer) with speedup of 6.25×6.25$\times$6.25 ×, 1.15×1.15$\times$1.15 ×, and 1.75×1.75$\times$1.75 × in total latency for the training workload (JOB, TPC-DS, Stack) and 8.33×8.$3$, 1.15×1.15$\times$1.15 ×, and 1.25×1.25$\times$1.25 × for the testing workload, respectively. Compared to other learned optimizers, FOSS consistently outperforms the other four methods in terms of average performance for both training and testing workloads. FOSS outperforms Bao, HybridQO, Loger and Balsa with average speedup of 1.74×1.74$\times$1.74 ×, 2.04×2.04$\times$2.04 ×, 1.21×1.21$\times$1.21 × and 1.90×1.90$\times$1.90 × on training workloads, and 1.49×1.49$\times$1.49 ×, 1.67×1.67$\times$1.67 ×, 1.33×1.$3$ and 2.28×2.28$\times$2.28 × on testing workloads, respectively.
Specially, the JOB workload entails complex query statements and non-uniform data distributions, posing a significant challenge to query optimizers. Therefore, the performance on the JOB workload is more indicative of an optimizer’s capabilities. FOSS significantly outperforms other SOTA methods on the JOB workload, demonstrating its superior performance.
As shown in Table I, FOSS not only outperforms other methods in terms of total workload latency but also demonstrates superior performance in the GMRL metric. Compared to Balsa, FOSS benefits from the assurance provided by the original plans and efficiently interacts with the simulated environment, enabling more exploration. In contrast to Loger, Bao, and HybridQO, FOSS utilizes finer-grained optimization, increasing the likelihood of generating superior plans than original plans. These factors contribute to FOSS outperforming other methods in terms of the GMRL metric and surpassing the expert optimizer across a wider range of queries, highlighting its stronger optimization capabilities. This will be further confirmed in VI-B4.
VI-B2 Training efficiency
Fig. 5 illustrates the training curves of the average speedup relative to the expert optimizer over time for different learned optimizers on the test workload. The shaded area indicates the range between the minimum and maximum values. It can be observed that benefiting from the assurance provided by the original plans, FOSS can swiftly demonstrate its optimization effects and surpass the performance of the expert optimizer. On the JOB workload, FOSS achieves superior performance compared to other SOTA methods after 5 hours of training and converges to the best performance level after around 8 hours. On the TPC-DS workload, FOSS converges in about 1 hour and outperforms Loger, Balsa, and HybridQO, reaching performance levels comparable to Bao. On the Stack workload, FOSS also converges in about 1 hour, exhibiting a training efficiency curve similar to Loger.
VI-B3 Optimization Time